PHOTOCHEMICAL STUDIES OF CH_{3}OH ISOLATED IN SOLID PARAHYDROGEN

\underline{DAVID T. ANDERSON}, KYLIE A. KUFELD, DOMINIQUE SCHOECH, ROBERT B. SLIPP, AND FARAJ M. ALMARRI, Department of Chemistry, University of Wyoming, Laramie, WY 82071-3838.

Our group is studying the 193.3 nm photochemistry of methanol (CH_{3}OH) isolated in solid pH_{2} using high-resolution FTIR spectroscopy to explore the low temperature reactions of H-atoms with CH_{3}OH. Gas phase studies have shown that the dominant photodissociation channel is CH_{3}OH + hν → CH_{3}O + H,^{a} and thus the in situ photochemistry is expected to produce methoxy and H-atom radicals. After photolysis the methoxy radicals are immobilized at these temperatures, but the H-atoms can still readily move through solid pH_{2}. For short time 193.3 nm exposures, greater than 90% of the CH_{3}OH remains intact and therefore repeated FTIR spectra immediately after photolysis can be used to record the kinetics of reactions of H-atoms with CH_{3}OH. We observe exponential growth in the intensities of a number of peaks that can be assigned to the hydroxymethyl CH_{2}OH radical. Further, the growth is bi-exponential and we believe there are two processes that lead to the formation of CH_{2}OH: (1) fast growth by interconversion of CH_{3}O to CH_{2}OH and (2) slow growth due to hydrogen abstraction reactions of cold H-atoms with CH_{3}OH. We have not observed the methoxy radical. Currently we are performing photochemical experiments on fully deuterated methanol (CD_{3}OD) in solid pH_{2} to test these kinetic interpretations, and the most recent results and analysis will be presented.

aY. Wen, J. Segall, M. Dulligan, and C. Wittig, J. Chem. Phys. 101, 5665 (1994).